PureSoulless:

Add all the numbers formed by squaring each of the three successive natural numbers and find the remainder when it is divided by 3.

adrianluvvsyouu2:

I solved this using Python, but you can definitely do this mathematically n^2 + (n+1)^2 + (n+2)^2 can be expanded into 3n^2 + 6n + 5 I hope you know of the mod function, which tells us the remainder a mod b = c This means a divided by b gives us a remainder of c Lets take a look at each term: 3n^2 mod 3 = 0, since 3n^2 is definitely divisible by 3 6n mod 3 = 0, since 6n is definitely divisible by 3 5 mod 3 = 2, since 3 only goes into 5 one time, and then there is a remainder of 2 So lets write that out as 3n^2 mod 3 + 6n mod 3 + 5 mod 3 = 0 + 0 + 2 = 2 The remainder then, should always be equal to 2! Hope this makes sense, and if you have questions I'm happy to answer

PureSoulless:

@adrianluvvsyouu2 wrote:
I solved this using Python, but you can definitely do this mathematically n^2 + (n+1)^2 + (n+2)^2 can be expanded into 3n^2 + 6n + 5 I hope you know of the mod function, which tells us the remainder a mod b = c This means a divided by b gives us a remainder of c Lets take a look at each term: 3n^2 mod 3 = 0, since 3n^2 is definitely divisible by 3 6n mod 3 = 0, since 6n is definitely divisible by 3 5 mod 3 = 2, since 3 only goes into 5 one time, and then there is a remainder of 2 So lets write that out as 3n^2 mod 3 + 6n mod 3 + 5 mod 3 = 0 + 0 + 2 = 2 The remainder then, should always be equal to 2! Hope this makes sense, and if you have questions I'm happy to answer
Do you know how to do it mathematically? also I'm supposed to explain it to a grade 9 student as a project, and I don't remember using modulo in grade 9.

PureSoulless:

Thank you though.

PureSoulless:

@silvernight269 @noodlesandriceyt